A car accelerates uniformly from rest and reaches an angular speed of 22.0 m/s in
9.00 s. Assuming the diameter of a tire is 58.0 cm, (a) find the number of revolutions the tire makes during this motion, assuming that no slipping occurs. (b) What is the final angular speed of a tire in revolutions per second?
1
Expert's answer
2020-10-27T11:11:19-0400

Given:

"\omega_0=0\: \rm rad/s"

"v=22.0\: \rm m/s"

"t=9.00\: \rm s"

"d=58.0\:\rm cm"

The final angular speed

"\omega=\frac{v}{R}=\frac{22.0}{0.58/2}=75.9\:\rm rad/s"

(a)

"\phi=\frac{\omega+\omega_0}{2}=\rm \frac{75.9+0.00}{2}\times 9.00=341\: rad"

"N=\frac{\phi}{2\pi}=\frac{341}{2\pi}=54.4"

(b)

"n=2\pi\omega=6.28\times 75.9\:\rm rad/s=477 \:\rm rev/s"